Derive an equation x(t) for the instantaneous position of the car as a function of time. Identify the value x when t=0s.use the given data for the time to reach 400m to calculate a constant of interaction c
Formula:
V(t) =A(1-e^-t/tmax)
Numbers =
t(400m) (s) =10.5
Tmax=7.1s
1
Expert's answer
2020-06-30T15:26:36-0400

"v=A(1-e^\frac{t}{tmax})"

"x=\smallint vdx=A\smallint 1-e\frac{-t}{tmax}dt"

"=A[t+tmax×e^\frac{-t}{tmax}]+C"

"400=A[10.5+7.1e^\frac{-10.5}{7.1}]+C"

"C=400-A(12.12)"


"X=A[t+tmax×e^\frac{-t}{tmax}]+400-12.12A"

"X=A[t+tmax×e^\frac{-t}{tmax}-12.12]+400"

At t=0

"X=A[0+7.1-12.12]+400"

"=400-5.02A"

"=394.98A" is the answer.